Problem 12.20

Premier Corp. has net sales of $1,000,000, and cost of goods sold equal to 76 percent of net sales. Assume all sales are credit sales. If the firm’s accounts receivable total $127,000 and its operating cycle is 79.0 days, how much inventory does the firm have? (Round answer to nearest dollar, e.g. 5,275.)

The firm’s inventory is $..............
Answer #1

account receivable days are

=365/(1000000/127000)

=46.355 days

inventory days=Operating cycle-account receivable days

=79-46.355

=32.645 days

how much inventory does the firm have

=(32.645/365)*(1000000*76%)

=67973

the above is answer..